OpenStudy (anonymous):

EASY AGAIN. solve 4sinx=cosec x . how ??? help me

OpenStudy (gorv):

@Ika_Kim 4sinx=1/sinx 4(sinx)^2=1 sinx=(1/4)^1/2 sinx=+-1/2 remember sinx not equal to 0

OpenStudy (anonymous):

ohh,,, we can just use square root

OpenStudy (anonymous):

okayyy thankss ^_^ i got it

OpenStudy (gorv):

@Ika_Kim yes plzz click best response if u understood it